OpenStudy (anonymous):

In making punch for the picnic, the ratio of ginger ale to fruit juice was 3 to 4. If 68 ounces of fruit juice was used, how many ounces of ginger ale was used?

OpenStudy (anonymous):

3/4=?/68 then cross multiply so (3x68/4) so the ounces of ginger ale should be 51

OpenStudy (anonymous):

\[\frac{3}{4}=\frac{x}{68} \]x = 51

OpenStudy (anonymous):

thanks math is my worst subject

OpenStudy (anonymous):

lol same here

OpenStudy (anonymous):

i just have a really amazing teacher but i still dislike math haha

OpenStudy (anonymous):

i have another question maybe you can help by showing me how to do it

OpenStudy (anonymous):

okay I will definitely try to help

OpenStudy (anonymous):

If ∠A and ∠B are supplementary angles and if m∠A is 68°, what is m∠B?

OpenStudy (anonymous):

okay well supplementary means 180 so that means angles a+B have to equal 180 so if a=68 you should put that in the problem and solve so a+b=180 68+b=180 so subtract 68 from both sides and you get b=112

OpenStudy (anonymous):

wow i dont know if i got smarter or more confused :)

OpenStudy (anonymous):

okay I don't know if this will help you or not but I know it's helpful to me so I'll tell you how i remember the difference between complimentary and supplementary(: so when i hear complimentary, i think of compliments right? and an acute( A CUTE) angle is less than 90 degrees. so like A CUTE angle like the tiny CUTE adorable angles are the 90 degrees and the supplementary ones are double because they're supplements haha idk just something i use

OpenStudy (anonymous):

and i can try to draw it for you if you'd like?

OpenStudy (anonymous):

no thanks i would really be confused then i will study it more thanks for the help

OpenStudy (anonymous):

okay sorry i feel like i confused you more :3

OpenStudy (anonymous):